### v1

Abstract
The current evaluation of mathematical skills in LLMs is limited, as existing benchmarks are either relatively small, primarily focus on elementary and high-school problems, or lack diversity in topics. Additionally, the inclusion of visual elements in tasks remains largely under-explored.
To address these gaps, we introduce U-MATH, a novel benchmark of 1,100 unpublished open-ended university-level problems sourced from teaching materials. It is balanced across six core subjects, with 20% of multimodal problems. Given the open-ended nature of U-MATH problems, we employ an LLM to judge the correctness of generated solutions. To this end, we release -MATH, a dataset to evaluate the LLMs’ capabilities in judging solutions.
The evaluation of general domain, math-specific, and multimodal LLMs highlights the challenges presented by U-MATH. Our findings reveal that LLMs achieve a maximum accuracy of only 63% on text-based tasks, with even lower 45% on visual problems. The solution assessment proves challenging for LLMs, with the best LLM judge having an F1-score of 80% on -MATH
We open-source U-MATH, -MATH, and evaluation code on GitHub.111https://github.com/toloka/u-math
1 Introduction
Mathematical reasoning is a fundamental domain for assessing the true capabilities of Large Language Models (LLMs) to reason (Ahn et al., 2024). While existing benchmarks like GSM8K (Cobbe et al., 2021) and MATH (Hendrycks et al., 2021) provide valuable insights, they primarily focus on school-level mathematics. This leaves a significant gap in understanding how LLMs perform on more advanced, university-level problems. Moreover, these benchmarks are becoming saturated, as GPT-4, using advanced prompting techniques, has achieved over 92% success rate on GSM8K and 80% on MATH (Achiam et al., 2023).
Recent works, such as CHAMP (Mao et al., 2024) and MathOdyssey (Fang et al., 2024), aim to introduce more challenging problems but are limited in size (<400 samples) and lack comprehensive topic coverage. The most challenging problems stem from school-level competitions or olympiads, missing the crucial middle ground of university-level coursework that reflects academic demands.
Furthermore, there is a growing interest in assessing multi-modal LLMs’ abilities to perform mathematical reasoning involving visual elements (Ahn et al., 2024). Large datasets like MathVista (Lu et al., 2023), We-Math (Qiao et al., 2024), or MathVerse (Zhang et al., 2024) provide an extensive set of (mostly) visual tasks but may lack university-level problems and often rely on multiple-choice validation, leading to easier problems and faster saturation of benchmarks.
In turn, evaluating complex free-form answers remains a significant challenge for the field (Hendrycks et al., 2021). Current methods often rely on LLM judges to assess problems, which introduces potential biases and inconsistencies (Zheng et al., 2023). Errors introduced by automatic evaluators are often overlooked in popular benchmarks. This oversight makes it impossible to account for judge biases, which detracts from the reliability of the evaluation results.
Recent studies also indicate that evaluation of mathematical solutions is a demanding task (Zeng et al., 2023; Xia et al., 2024) and that an LLM’s ability to judge mathematical solutions is correlated with its problem-solving performance (Stephan et al., 2024), further signifying the importance of evaluations designed to asses the evaluators themselves — also called meta-evaluations.
Popular datasets for the task of mathematical meta-evaluation are PRM800K (Lightman et al., 2023), MR-GSM8K (Zeng et al., 2023) and MR-MATH (Xia et al., 2024). However, these are all based on the GSM8K and MATH datasets, still leaving a gap in meta-evaluations for university-level problems.
Aiming to bridge these gaps and provide a comprehensive evaluation of LLMs’ mathematical capabilities, we introduce U-MATH (University Math) and a supplementary meta-evaluation dataset, which we refer to as -MATH (Meta U-MATH). Our main contributions are:
1.
U-MATH Benchmark (Section 3): We open-source a set of 1,100 of university-level problems collected from actual coursework with final answers and solutions. About 20% of problems require image understanding to be solved. The text-only part of the benchmark is balanced across 6 key subjects: Precalculus, Algebra, Differential Calculus, Integral Calculus, Multivariable Calculus, and Sequences&Series.
2.
-MATH Meta-Evaluation Benchmark (Section 3.3): Additionally, we introduce a set of 1084 meta-evaluation tasks sourced from U-MATH problems and designed to rigorously assess the quality of LLM judges. We manually select approximately 25% of the U-MATH problem statements and golden answers, supplying each with four solutions produced by different top-performing LLMs, and label them based on whether the generated solutions are correct or not. The benchmark is designed to be challenging for LLM judges yet representative of the typical university-level math grading tasks.
3.
Comparison of Models (Section 4): We conduct a comparative analysis of various open-source and proprietary LLMs on U-MATH. Our analysis highlights the high performance of specialized models in text-only problems and the superiority of proprietary models in visual tasks with the best U-MATH accuracy of 49%. Additionally, we examine several popular LLMs on -MATH to assess their ability to judge free-form mathematical problems. Our results show the best model achieving the macro F1-score of 80%.
We release the U-MATH and -MATH benchmarks under a permissive license to facilitate further research and ensure reproducibility.
2 Background
Enhancing and evaluating the mathematical reasoning capabilities of LLMs is essential in AI research (Ahn et al., 2024). Studies show that finetuning with mathematical and code-related data enhances models’ general skills (Prakash et al., 2024). Mathematical tasks require logical thinking and multi-step problem-solving, thus improving overall reasoning abilities in LLMs (Chen et al., 2024).
This leads to the problem of evaluating LLM’s math abilities. Despite the significant progress, many existing benchmarks are limited in scope, focusing primarily on school-level mathematics or limited in size and topic coverage. Table 1 summarizes popular text-only and visual mathematical benchmarks.
Dataset Levels %Uni. Level #Test %Visual %Free Form Answer MMLU (Hendrycks et al., 2020) 0 1.3k 0 0 GSM8k (Cobbe et al., 2021) 0 1k 0 0 MATH (Hendrycks et al., 2021) 0 5k 0 100 MiniF2F (Zheng et al., 2021) 0 244 0 100 OCWCourses (Lewkowycz et al., 2022) 100 272 0 100 ProofNet (Azerbayev et al., 2023) 50 371 0 100 CHAMP (Mao et al., 2024) 0 270 0 100 MathOdyssey (Fang et al., 2024) 26 387 0 100 MMMU (Yue et al., 2023) 0 505 100 0 MathVista (Lu et al., 2023) 0 5k 100 46 MATH-V (Wang et al., 2024a) 0 3k 100 50 We-Math (Qiao et al., 2024) 20 1.7k 100 0 MathVerse (Zhang et al., 2024) 0 4.7k 83.3 45 U-MATH (this work) 100 1.1k 20 100
Textual Mathematical Benchmarks.
Early efforts to assess LLMs’ mathematical abilities have emerged in datasets like MathQA (Amini et al., 2019) and the mathematics subset of MMLU (Hendrycks et al., 2020). These early benchmarks emphasized the importance of operation-based reasoning in solving mathematical word problems, typically in a multiple-choice format. Nowadays, even smaller models (e.g., 7B parameters) have achieved high scores on these tasks (Li et al., 2024b), suggesting that these benchmarks are becoming saturated. In response, more comprehensive datasets have emerged, such as GSM8K (Cobbe et al., 2021) and MATH (Hendrycks et al., 2021), or MGSM (Shi et al., 2022) (multilingual version of 250 GSM8K samples). These popular benchmarks are crucial for evaluating LLMs’ mathematical reasoning skills. However, they primarily focus on school-level problems, which may not fully assess the depth of mathematical reasoning.
Recent efforts attempt to address more advanced mathematical concepts. MathOdyssey (Fang et al., 2024) with competition problems, OCWCourses (Lewkowycz et al., 2022) from actual MIT courses, and ProofNet (Azerbayev et al., 2023) focusing on proofs aim to evaluate undergraduate-level or olympiad-level knowledge. However, these datasets are constrained by their small sizes (e.g., 387, 272, and 371 samples), limiting their statistical robustness and topic coverage. For example, MathOdyssey is limited to 101 samples in university-level topics (Calculus, Algebra, and Diff. Equations and Statistics). Other specialized datasets like MiniF2F (Zheng et al., 2021) provide valuable parallel corpora in formal languages, while CHAMP (Mao et al., 2024) offers helpful context and hints, but both are similarly limited in scale with 244 and 270 samples. Additionally, both heavily rely on already published resources: CHAMP sources material from a book, while MiniF2F re-uses international olympiads and MATH dataset problems. An attempt to provide a more robust evaluation, GHOSTS (Frieder et al., 2024) dataset, provides 728 problems (both from other datasets and new ones) but does not provide reference solutions and answers, focusing instead on human evaluation, making cheap automatic evaluation impossible.
The current datasets are either too small, leading to higher measurement errors, or focus mainly on elementary and high school math, leaving a gap in evaluating LLMs’ proficiency in advanced university-level math topics.
Visual Mathematical Benchmarks.
As multimodal LLMs gain prominence, there is a growing need for visual mathematical benchmarks (Zhang et al., 2024; Qiao et al., 2024). Early efforts in this domain focus primarily on geometric problems, as seen in datasets like GeoQA (Chen et al., 2022b), UniGeo (Chen et al., 2022a), and Geometry3K (Lu et al., 2021). These datasets have a narrow focus that does not encompass the breadth of mathematical visual reasoning required at advanced levels.
More recent benchmarks attempt to broaden the scope of visual mathematical evaluation. One of the first comprehensive attempts is the mathematical subset of MMMU (Yue et al., 2023), which offers 505 college-level multiple-choice questions, all with images. However, its multiple-choice format limits the complexity of problems that can be posed. MathVista (Lu et al., 2023) collects 28 existing datasets and introduces 3 new datasets with a total of 5k samples (1k testmini samples). However, as shown by Qiao et al. (2024), it faces challenges with data quality due to its compilation from older datasets.
The latest benchmarks, such as MATH-V (Vision) (Wang et al., 2024a) and We-Math (Qiao et al., 2024), extend this approach to collect 3k and 1.7k visual samples, respectively. However, both datasets rely on multiple-choice questions in the test set, leading to faster saturation. MathVerse (Zhang et al., 2024) further extends this approach, relying on visual elements and providing some simple text problems with 1.2k brand-new samples. Among these, only the We-Math dataset includes university-level mathematical problems.
Our U-MATH dataset improves on existing benchmarks with 225 of 1,100 university-level problems that require visual elements (graph, table, diagram) to be solved. This balanced ratio ensures models are challenged to handle both traditional and visual problem-solving without over-relying on visuals, mirroring real-world scenarios.
Large Language Models for Mathematics.
The application of LLMs to mathematical problem-solving shows promising results, particularly with models like GPT-3.5 and GPT-4 demonstrating strong reasoning abilities for complex tasks such as those in the MATH dataset (Achiam et al., 2023). While open-source models initially lagged in performance on advanced mathematical tasks, the Llama-3.1 (Dubey et al., 2024) is approaching parity with proprietary models. The most popular benchmarks, MATH and GSM8K, are nearing saturation, with Llama 3.1 405B achieving scores of 73.8% and 96.8%, respectively. Similarly, a Qwen2.5-Math-72B model (Yang et al., 2024b; Team, 2024) reach 85.9% on MATH while Qwen2-Math-72B (Yang et al., 2024a) reaches 96.7% on GSM8k.
To enhance LLMs’ mathematical capabilities, researchers develop various prompt-based methods (Liu et al., 2021). These include techniques for encouraging chain-of-thought generation (Wei et al., 2022), selecting final results from multiple sampled outputs (Wang et al., 2022), and using external tools such as calculators, WolframAlpha or Python interpreters (Gao et al., 2023) to reduce arithmetic errors. Additionally, instruction tuning during pre-training has been identified as a key factor in improving performance (Wang et al., 2017). While these approaches show promise, their effectiveness on university-level problems still needs to be explored due to the lack of suitable large-scale benchmarks.
Mathematical solution verification.
Evaluating mathematical solutions is uniquely challenging due to the open-ended nature of answers and the inherent ambiguity in mathematical expressions. Consequently, many benchmarks opt for multiple-choice formats due to their grading simplicity. However, this approach often simplifies tasks, providing hints that models can exploit (Li et al., 2024c; Pezeshkpour and Hruschka, 2023).
While free-form evaluation using LLM judges is widespread (Zheng et al., 2023), it is known to introduce potential errors (Zheng et al., 2023), since evaluating mathematical solutions is a complex task in its own right (Zeng et al., 2023; Xia et al., 2024). These evaluation errors are largely overlooked and unaccounted for, limiting the reliability of inferences drawn from such evaluations.
Hence, it is important to be able to estimate the performance of automatic evaluators and to choose the most adequate among them. Recent studies show that evaluation performance is correlated with but does not equal problem-solving performance (Stephan et al., 2024). This underscores the importance of benchmarks designed specifically to asses the evaluators — also called meta-evaluations.
There are existing benchmarks that are well-suited for meta-evaluations. PRM800K (Lightman et al., 2023) contains 800K annotated steps from 75K solutions to 12K MATH dataset problems, designed to confuse reward models. FELM (Zhao et al., 2024) provides GPT-3.5 annotations for solutions to 208 GSM8K and 194 MATH problems. MR-GSM8K (Zeng et al., 2023) and MR-MATH (Xia et al., 2024) introduce meta-evaluation datasets focused on the GSM8K and MATH datasets, respectively. However, these are either based on elementary to high-school level problems or feature specifically competition-style math, leaving a gap in meta-evaluations on complex and practical university tasks.
To address this, we introduce -MATH— a meta-evaluation dataset based on a subset of U-MATH problems. It provides LLM-generated solutions with verified labels, enabling precise and fine-grained assessment of LLMs’ evaluation abilities.
3 U-MATH
We present U-MATH (stands for University Math) — a benchmark designed to challenge LLMs with problems requiring deep understanding and advanced reasoning. The problems span 6 core topics and range in difficulty and number of questions. A subset of 20% of problems includes images to test the models’ ability to interpret and reason with graphical information. Reference solutions and answers accompany all problems.
Accuracy is the primary performance metric for U-MATH, its text-only problems (U-MATH) and problems that include a visual component (U-MATH). The main performance measure for -MATH is macro-F1.
We use an LLM as a judge (Zheng et al., 2023) to measure the accuracy of the free-form answers against the golden solutions. A problem is considered solved only if all required questions are answered and all requested items (e.g., all saddle points) are correctly identified.
3.1 Dataset Collection
To create a benchmark that authentically reflects university-level mathematics, we collaborate with Gradarius, a platform providing learning content and software for top US universities specialized in mathematics. The problems are sourced from ongoing courses across various institutions currently run on the Gradarius platform. Problems and solutions are crafted by subject matter experts and represent real-world academic standards. These samples are unpublished and have not been exposed to any external sources. Thus, the dataset could not be leaked to current LLMs.
We employ a multi-stage filtering process to select challenging problems from tens of thousands of available samples. First, we filter out problems with short solutions ( characters) and problems in multiple-choice format. As LLMs are not designed to perform arithmetic calculations and are prone to errors (Hendrycks et al., 2021; Lewkowycz et al., 2022), we focus on testing mathematical reasoning rather than calculations. We filter out problems marked as allowing calculator usage. As for the visual problems selection, we chose to keep problems with a single image for convenience.
Next, we employ several small LLMs (LLaMA-3.1-8B (Dubey et al., 2024), Qwen2-7B (Yang et al., 2024a), Mistral-7B (Jiang et al., 2023), Mathstral-7B, NuminaMath-7B (Beeching et al., 2024)) to solve the problems. We select 150 most challenging problems for each subject based on the average problem solution rate. For this step, we use the same pipeline as described in Section4. This way, we ensure that none of the individual models influence problem selection largely and that there is no overfitting to a specific LLM. As the last step, we hold extra validation high risk problems (with low solve rate) using our in-house math experts and Gradarius content team.
After collection, we enlist a team of experts from the Stevens Institute of Technology, who actively teach various Calculus courses. The experts verify that each problem is suitable either for assessing the subject knowledge expected of college or university students or for testing prerequisite knowledge.
The team thoroughly reviewed and affirmed that the selected problems meet these criteria.
Overall, only 4.3% of the problems are categorized as school-level rather than university-level, highlighting the robustness of the selection process.
3.2 Dataset Statistics
The U-MATH benchmark comprises 1,100 carefully curated and validated mathematical problems. These problems are distributed across 6 core subjects with about 20% of the tasks incorporating visual elements, such as graphs, tables, and geometric figures, mirroring the multi-modal nature of real-world mathematical problems: Precalculus (Review), Algebra, Differential Calculus (+Differential Equations), Integral Calculus, Multivariable Calculus, and Sequences & Series.
Math Subject #Textual #Visual Avg. Questions Avg. Answers Algebra 150 30 1.93 1.28 Differential Calculus 150 70 2.37 1.15 Integral Calculus 150 58 1.09 1.01 Multivariable Calculus 150 28 1.74 1.09 Precalculus 150 10 1.51 1.23 Sequences and Series 150 4 1.36 1.00 All 900 200 1.66 1.12
Table 2 summarizes the distribution of problems across different subjects. The average is 1.7 questions per problem (e.g., local minima, maxima, and increasing intervals are asked), and the average of 1.1 answers per question (for example, the number of saddle points in the correct answer).
3.3 Meta-Evaluation Framework (-MATH)
The evaluation of mathematical problems is not straightforward. Even simple expressions such as may have valid forms like , , , or unsimplified variants like . In practice, evaluating free-form solutions requires testing expression equivalence in much less trivial cases, especially with more advanced problems (refer to Section A.3 in Appendix for an example).
To systematically study the ability of LLMs to evaluate free-form mathematical solutions on advanced, university-level problems, we introduce the -MATH (Meta U-MATH) benchmark. It consists of a curated subset of U-MATH samples, supplied with LLM-generated solutions — both correct and not. The solutions are labeled using a combination of manual inspection and automated verification via Gradarius-API, which allows to test formal equivalence of mathematical expressions.
We selected 271 U-MATH problems (around 25%) based on their assessment difficulty to create a challenging meta-evaluation set. This subset does not aim to reflect the overall U-MATH distribution but rather to provide a robust test for LLM judges. We focused on text-only problems, excluding those needing images, due to the limited size of the labeled U-MATH subset. Four solutions have been generated for each of the selected problems — using Qwen2.5-72B, Llama3.1-8B, GPT-4o and Gemini-1.5-Pro models — 1084 samples in total.
A tested model is provided with a problem statement, a reference answer, and a solution to evaluate. We treat this as a binary classification task, using the macro-averaged F1-score as the primary metric to minimize the effect of class imbalance. Additionally, we report Positive Predictive Value (PPV or Precision) and True Positive Rate (TPR or Recall) for the positive class as well as Negative Predictive Value (NPV) and True Negative Rate (TNR) for the negative class, offering a finer-grained performance evaluation. We also report all of the scores computed both across the entire set of samples and only across those with solutions produced by a specific model, separately for each of the author models.
4 Experiments and Results
4.1 Experimental Setup
We select some top-performing recent LLMs to evaluate.
Model Source Size(s) Visual Open-weights Mathstral-v0.1 (Mistral.ai, 2024) 7B NuminaMath-CoT (Beeching et al., 2024) 7B LLaMA-3.1 (Dubey et al., 2024) 8B, 70B LLaMA-3.1-Nemotron (Wang et al., 2024b) 70B Qwen2-Math (Yang et al., 2024a) 7B, 72B Qwen2.5-Math (Yang et al., 2024b) 7B, 72B Qwen2.5 (Team, 2024) 7B, 72B Athene-V2-Chat (Nexusflow, 2024) 72B Pixtral-12B-2409 (Mistral AI, 2024) 12B LLAVA One Vision Qwen2-7B (Li et al., 2024a) 8B Qwen2-VL (Yang et al., 2024a) 7B, 72B LLaMA-3.2 (Meta AI, 2024) 11B, 90B Claude-3.5-Sonnet (Anthropic, 2024) unknown GPT-4o-mini-2024-07-18 (OpenAI, 2024) unknown GPT-4o-2024-08-06 (OpenAI, 2024) unknown Gemini-1.5-Flash-002 (Team et al., 2024) unknown Gemini-1.5-Pro-002 (Team et al., 2024) unknown
All LLMs are tested using the same prompts and settings for fair comparison. The LLMs are restricted to a single generation of 4096 tokens with the temperature set to 0. We employ chain-of-thought (CoT) prompting (Wei et al., 2022) to encourage models to ‘think’ before providing an answer. mages are included directly in the prompts for multimodal LLMs. To text-only LLMs the problem description is provided as-is without visual elements.
We report accuracy based on GPT-4o-2024-08-06 as-a-judge for our final results, despite it not being the best performing judge — due to the model still residing among the top-ranked judges, being the more conservative one in terms of false positive rate, as well as widely available, leading to easier reproducibility. Details on the judgment setup and comparisons of various judges are all discussed in Section 4.3.
4.2 U-MATH Results
Figure 2 compare popular text-only and multimodal models in U-MATH as well as U-MATH and U-MATH. Table 4 summarizes the performance of all evaluated LLMs on the U-MATH benchmark. Reference to Appendix E for model performance vs model size comparison.
Model U-MATH U-MATH Algebra Diff. C. Integral C. Multivar C. Precalculus Seq.& Series T V T V T V T V T V* T V* T V* 900 200 150 30 150 70 150 58 150 28 150 10 150 4 Text-only models Mathstral 7B 18.0 20.7 6.0 51.3 6.7 4.0 10.0 1.3 1.7 8.0 3.6 48.7 10.0 10.7 0.0 NuminaMath 7B 19.2 22.8 3.0 62.7 0.0 4.0 7.1 1.3 0.0 6.0 3.6 51.3 0.0 11.3 0.0 Llama-3.1 8B 22.3 26.1 5.0 59.3 3.3 6.7 5.7 9.3 3.4 11.3 3.6 54.7 10.0 15.3 25.0 Qwen2.5 7B 33.8 40.0 6.0 86.0 10.0 12.7 1.4 10.0 12.1 26.7 3.6 75.3 0.0 29.3 0.0 Qwen2.5-Math 7B 38.4 45.2 7.5 87.3 6.7 18.7 5.7 8.0 10.3 36.0 10.7 80.7 0.0 40.7 0.0 NuminaMath 72B 25.0 29.7 4.0 74.7 3.3 6.7 4.3 4.0 3.4 11.3 3.6 62.7 10.0 18.7 0.0 Llama-3.1 70B 28.5 33.7 5.0 82.0 3.3 10.7 5.7 4.0 5.2 14.0 3.6 64.0 0.0 27.3 25.0 Llama-3.1 Nemotron 70B 31.4 37.4 4.0 84.0 0.0 14.7 2.9 4.0 3.4 25.3 7.1 64.0 20.0 32.7 0.0 Qwen2.5 72B 41.0 48.6 7.0 88.7 6.7 22.7 4.3 12.0 6.9 40.0 17.9 83.3 0.0 44.7 0.0 Athene-V2 72B Chat 46.2 54.6 8.5 88.7 3.3 34.0 4.3 16.0 6.9 50.7 21.4 88.7 10.0 49.3 50.0 Qwen2.5-Math 72B 50.2 59.0 10.5 92.7 6.7 35.3 7.1 20.7 17.2 58.0 7.1 90.0 0.0 57.3 50.0 Multimodal models Pixtral 12B 15.5 15.6 15.5 44.7 23.3 1.3 34.3 0.7 0.0 3.3 0.0 32.0 0.0 11.3 0.0 Llama-3.2 11B Vision 17.0 18.6 10.0 54.0 10.0 1.3 20.0 1.3 1.7 4.7 3.6 43.3 10.0 6.7 0.0 LLaVA-OV Qwen2-7B 17.7 20.7 4.5 60.7 6.7 4.0 5.7 1.3 1.7 5.3 3.6 43.3 10.0 9.3 0.0 Qwen2-VL 7B 20.4 21.4 15.5 62.7 10.0 4.7 32.9 0.7 5.2 6.7 7.1 45.3 0.0 8.7 0.0 Qwen2-VL 72B 31.2 32.2 26.5 80.7 26.7 9.3 40.0 2.0 13.8 14.7 28.6 65.3 10.0 21.3 0.0 Llama-3.2 90B Vision 32.6 36.3 16.0 85.3 26.7 10.7 25.7 2.7 1.7 22.7 7.1 65.3 20.0 31.3 25.0 Claude Sonnet 3.5 35.1 36.1 30.5 76.0 33.3 12.0 41.4 7.3 17.2 21.3 28.6 65.3 30.0 34.7 25.0 GPT-4o-mini 37.2 40.3 23.0 88.0 16.7 16.7 31.4 4.0 10.3 24.0 35.7 77.3 20.0 32.0 25.0 GPT-4o 43.5 46.4 30.0 91.3 30.0 18.7 32.9 10.0 20.7 41.3 42.9 79.3 30.0 38.0 25.0 Gemini 1.5 Flash 51.3 53.8 40.0 91.3 50.0 36.0 45.7 14.0 24.1 44.0 50.0 80.7 30.0 56.7 50.0 Gemini 1.5 Pro 60.1 63.4 45.0 91.3 60.0 50.7 47.1 27.3 24.1 60.7 57.1 87.3 70.0 63.3 50.0
Among text-only models, the math-specific model Qwen2.5-Math-72B achieves the highest overall accuracy at 50.2%, showcasing strong mathematical reasoning capabilities. In the multi-modal model group, Gemini-1.5-pro-002 leads with an overall accuracy of 60.1%, highlighting the advantages of integrating visual processing. In contrast, best open-weights model Qwen2-VL-72B lacks mathematical abilities in visual and textual tasks with 31.2% on a U-MATH benchmark. Building on these results, several key trends emerge:
Model Size vs. Specialization: Larger models expectedly outperform smaller ones. However, the small specialized model Qwen2.5-Math-7B surpasses or performs on par with 10 times larger models like Qwen2.5-72B or LLaMA-3.1-70B and almost reaching leading Gemeni-1.5-Pro level. On the other hand, Pixtral-12B performs consistently worse than minor Qwen2-VL-7B, indicating a lack of university-level data in training.
Textual vs. Visual Problem-Solving: Across multimodal models, text-only problems’ accuracy vastly exceeds visual problems, highlighting areas for further improvement. The text-only models can solve a small percentage of visual problems, primarily due to guessing or judgment errors discussed in Section 4.3.
Proprietary vs. Open-weights model: Proprietary models like Gemini still offer top or competitive performance but lack transparency and flexibility. At the moment, the gap is evident in visual comprehension, with 18.5% difference on U-MATH between top-1 and best open-weight model. However, open-weight models like Qwen-Math is a big step toward top performance.
Continuous Finetuning: Additional tuning significantly enhances performance, with LLaMA-3.1 70B LLaMA-3.1 Nemotron 70B and Qwen2.5-72B Athene-V2 72B achieving 2.9% and 5.2% higher U-MATH accuracy, respectively. This reinforces the idea that models are not fully optimized for their size and require high-quality data for further improvements.
Subject-Specific Results
Model performance varies across mathematical subjects, excelling in text-based tasks for Precalculus and Algebra, consistent with benchmark saturation (Ahn et al., 2024), but faltering on visual-symbolic tasks. In Sequences and Series, success on formula-based problems reflects logical structuring, though limited visual data restricts evaluation. Differential and Multivariable Calculus results are moderate, with difficulties in abstract, multi-dimensional concepts, especially visual interpretations. Integral Calculus presents the greatest challenge, as interpreting curves, areas, and extensive expressions confounds models, underscoring the need for improved multimodal training.
4.3 Meta-Evaluation (-MATH) Results
For meta-evaluation we use the same setup as described in Section 4.1. Additionally, we experiment with two distinct prompting schemes — a standard Automatic Chain-of-Thought (AutoCoT) prompt involving a simple task description together with an instruction to think step-by-step, and a manual Chain-of-Thought prompt (which we refer to as simply CoT) with explicit instructions on which steps to follow when approaching the task. We find the latter prompting scheme to perform best, so we use manual CoT for the main results. The judge’s output is also further processed by an extractor model (Qwen2.5 72B is fixed for consistency), prompted to produce a single label — either ‘Yes’, ‘No’ or ‘Inconclusive’. We include ‘Inconclusive’ for cases when judge refuses to evaluate or generation fails; such judgments are treated as incorrect. Reference Appendix C.2 for full contents of all the prompts.
Model U-MATH -MATH -MATH -MATH -MATH -MATH F1 / F1 TPR TNR PPV NPV F1 / F1 F1 / F1 F1 / F1 F1 / F1 Llama-3.1 8B 26.1 52.0 / 53.1 48.7 55.9 56.0 48.5 48.7 / 49.6 49.2 / 51.2 51.2 / 57.6 55.5 / 50.5 Qwen2.5 7B 40.0 69.3 / 67.0 78.7 59.8 69.3 70.8 62.4 / 60.5 72.3 / 72.4 68.3 / 66.4 69.1 / 65.0 Qwen2.5-Math 7B 45.2 61.9 / 61.2 76.6 47.9 62.9 63.9 59.7 / 56.7 63.8 / 64.0 57.2 / 58.5 63.8 / 61.2 GPT-4o-mini 40.3 72.3 / 69.2 59.0 88.1 85.1 65.1 69.3 / 61.7 76.2 / 78.5 70.4 / 69.8 69.6 / 64.3 Gemini 1.5 Flash 53.8 74.8 / 65.3 63.3 88.3 86.2 67.6 71.2 / 61.9 80.6 / 70.8 70.1 / 65.3 73.9 / 59.7 Llama-3.1-70B 33.7 61.0 / 68.2 62.5 59.6 64.1 57.9 56.0 / 63.8 57.0 / 70.2 69.4 / 69.8 58.8 / 64.4 Qwen2.5 72B 48.6 75.6 / 75.1 77.1 74.2 77.5 73.7 70.5 / 68.9 79.3 / 80.1 73.7 / 73.4 74.2 / 73.8 Qwen2.5-Math 72B 59.0 74.0 / 75.5 80.9 66.8 73.8 75.2 69.3 / 68.8 77.3 / 79.8 68.2 / 69.2 76.8 / 80.4 Claude 3.5 Sonnet 36.1 74.8 / 68.1 62.5 89.5 87.3 67.4 70.8 / 64.1 77.9 / 71.8 72.2 / 68.1 73.8 / 63.4 GPT-4o 46.4 77.4 / 74.2 70.1 85.9 85.1 71.3 74.2 / 68.2 81.8 / 78.9 77.5 / 75.8 72.6 / 70.5 Gemini 1.5 Pro 63.4 80.7 / 69.1 77.5 84.5 85.2 76.4 77.7 / 64.6 83.6 / 74.8 78.2 / 68.6 79.5 / 64.9
We find that using manual CoT instructions instead of the standard AutoCoT improves or maintains judgment performance, save for Llama models, as shown in Table 5. Llama’s performance drop is largely due to increased inconclusive judgment rates (see Appendix G). At the same time, Gemini models benefit the most from this transition, gaining over 10% in F1-score and becoming the top-ranked models, surpassing Qwen and GPT models that outperform Gemini in the AutoCoT setting. This shows that prompting effects are substantial yet inhomogeneous across models. Please refer to Appendix F for a visual comparison.
In terms of the resulting performance, we see that correctly identifying a positive label is harder on average compared to negative labels, with the best TPR being almost 10% lower than the best TNR, and that the best attainable F1 score is only 80.7%. This constitutes a considerable deficiency in the context of judgment, because judges’ error rates directly limit the precision of capability evaluations, potentially even biasing them in case the errors are systematic in nature as opposed to pure noise.
Our results, for instance, reveal a consistent bias towards some models — better performance on Llama solutions and worse performance on Qwen solutions — most pronounced with smaller-sized judges and AutoCoT prompting. This bias is generally reduced for both small and large judges when transitioning to CoT prompting, which is also illustrated with Figure 3. At the same time, no noticeable ‘self-judgment’ effects are found.
It is also evident that being a better solver does not necessarily lead to being a better judge. In fact, our results suggest a trade-off existing between these skill; refer to Appendix H for visualizations and a more detailed discussion.
Besides that, we observe substantive differences in judges’ behavior: proprietary models tend to be more conservative — having relatively high TNR compared to their TPR — while Qwen family of models exhibits the opposite pattern. The behavior differences are further studied and illustrated in Appendix I.
Overall, judges show varying behaviors, they have imperfect performance that is also distinct from problem solving performance, and different prompting schemes induce nontrivial changes in judges’ behaviors, biases, and even their performance rankings.
All of these findings underscore the importance of performing meta-evaluations, since such things are impossible to quantify and comparisons impossible to make in the absence of datasets designed to benchmark the judges.
5 Conclusion
We introduce U-MATH, a novel multimodal benchmark for evaluating the university-level mathematical reasoning of LLMs. U-MATH includes 1,100 unpublished free-form problems from real teaching materials, covering 6 core mathematical subjects, with 20% involving image-based reasoning. Additionally, we provide -MATH, a meta-evaluation dataset, to assesses LLMs’ ability to evaluate free-form mathematical solutions.
Our experiments highlight significant challenges for LLMs in advanced reasoning and visual problem-solving. The highest accuracy achieved was 63.4% on text-based tasks and 45.0% on visual problems (Gemini-1.5-pro-002). Solution assessment remains difficult, with Gemini hiy top -MATH F1-score of 80%, showing room for improvement and underscoring the limitations of widely used models like GPT-4o in evaluation tasks.
Limitations.
While U-MATH offers diverse university-level problems, it does not cover the full range of advanced topics and may introduce biases by favoring certain problem types. Also, selection process may introduce biases, potentially favoring certain problem types or difficulty levels (e.g., more accessible topics like Precalculus and Algebra). The inclusion of 20% visual problems, yet reflect real distribution, limits the evaluation of visual reasoning. Furthermore, reliance on LLMs for valuation introduces potential, as models struggle with complex reasoning and instructions, evidenced by our findings with the -MATH. The -MATH dataset encompass of 25% of U-MATH problems narrows the evaluation scope, but provide 4 diverse model families as solution generators.
Future Work.
Future research can focus on enhancing LLM performance by integrating existing tool-augmented models and exploring their effectiveness on U-MATH and -MATH tasks. For instance, incorporating external tools, such as formal solvers, could improve complex textual and multimodal reasoning capabilities. Additionally, our findings indicate that widely used models like GPT-4o are not a silver bullet for solution evaluation; thus, developing specialized (finetuned) models or techniques for more accurate and unbiased assessment is a promising direction. Expanding -MATH with formal verification methods could further enhance the evaluation processes. Additionally, conducting deeper prompt sensitivity analyses would provide valuable insights for the field.
By open-sourcing U-MATH, -MATH, and the evaluation code, we aim to facilitate further research in advancing the mathematical reasoning capabilities of LLMs and encourage the development of models better equipped to tackle complex, real-world mathematical problems.
Acknowledgement
We thank all contributors from Stevens Institute and Toloka.ai experts who assisted in sourcing and verifying problems, inspect the solutions and provided valuable feedback throughout the development of U-MATH.
We would like to give special thanks to the dedicated team of Gradarius experts who played a crucial role in validating problems and ensuring their quality: Jan Cannizzo Paul Schwartz Andrey Nikolaev Arina Voorhaar Chloe Weiers Funda Gul Tatiana Zenkevich Igor Teplukhovskiy Anastasiia Feklina Ruslan Akmetdinov Ekaterina Eremina Sofia Tekhazheva
Ethics Statement
We collected all data in U-MATH and -MATH with appropriate permissions, ensuring no personal or proprietary information is included. The datasets consist solely of mathematical problems and solutions, without any sensitive content. We open-sourced the datasets and code under suitable licenses to support transparency and research advancement. There are no known conflicts of interest associated with this work.
Reproducibility Statement
All datasets and code will be available on GitHub. Detailed descriptions of dataset collection and processing are in Section 3. The experimental setup, including model configurations and prompts, is outlined in Section 4, with full prompts provided in Appendices C.1 and C.2. These resources enable replication of our experiments.
Appendix A Problem examples
A.1 U-MATH Problems
A.2 U-MATH Problem and Solution
A.3 -MATH meta-evaluation
Appendix B Sub-topics distribution
The U-MATH dataset cover variety of topics across 6 core subjects. Below is the count of unique topics per subject:
Differential Calculus: 51 unique topics
Sequences and Series: 28 unique topics
Integral Calculus: 35 unique topics
Precalculus Review: 19 unique topics
Algebra: 74 unique topics
Multivariable Calculus: 53 unique topics
Subject Topic Count Topic Name Differential Calculus 29 Curve Sketching 13 Limits 12 One-Sided Limits 12 L’Hospital’s Rule 11 Increasing and Decreasing Functions 11 Higher Derivatives 10 Applications of Derivatives (Local Extrema) Sequences and Series 40 Taylor Series 30 Fourier Series 18 Maclaurin Series 12 Approximating Constants Using Power Series 6 Radius of Convergence (Center of Convergence) 5 Differentiate Power Series 4 Error in Approximation Integral Calculus 83 The Substitution Rule 24 Antiderivatives 10 Volumes of Solids of Revolution About the X-Axis 9 Trigonometric Substitutions and Inverse Substitutions 9 Integrate Respect Independent Variable 7 Applications of Integrals 7 Single Variable Surface Area Integrals Precalculus Review 55 Trigonometric Functions 24 Zeros 11 Inverses of Functions 8 Inequalities 7 Equations with Exponents and Logarithms 7 Properties of Functions 6 Exponential Functions Algebra 18 Equations and Inequalities 13 Polynomial Equations 8 Find Composition of Two Functions 7 Polynomials 6 Find Slope Line 6 Applications of Exponential Function 6 Quadratic Equations Multivariable Calculus 13 Triple Integrals 11 Lagrange Multipliers 9 Double Integrals in Polar Coordinates 8 Derivatives of Parametric Equations 8 Integrals of Multivariable Functions 8 Double Integral Over General Region 6 Classification of Critical Points
Appendix C Prompts
C.1 Prediction Prompt
C.2 Judgment Prompts
Appendix D Solution Predictions Length Distribution
Appendix E Model Accuracy vs Size
Appendix F -MATH Prompting Schemes Comparison
Appendix G -MATH Inconclusive Judgment Rates
Llama-3.1 8B Llama-3.1 70B Qwen2.5-Math 7B Qwen2.5-Math 72B Qwen2.5 7B Qwen2.5 72B GPT-4o-mini GPT-4o Gemini 1.5 Flash Gemini 1.5 Pro Claude 3.5 Sonnet AutoCoT 13.4 5.0 2.8 1.2 1.0 1.6 0.0 0.0 0.0 0.0 0.0 CoT 22.9 13.8 2.4 0.7 1.2 2.1 0.1 0.0 0.1 0.0 0.0
Appendix H Comparison of Problem Solving and Judgment Performance
In this section, we provide a detailed comparison of models’ performances on U-MATH and -MATH. The overall distribution of scores visualized in Figure 14 not only shows that improved problem-solving performance does not immediately lead to better judgment performance, as discussed in Section 4.3, but also suggests a possible trade-off present between these capabilities.
This possibility is further illustrated when considering specific models. For instance, Qwen2.5-Math demonstrates strong problem solving compared to most of the other models, but does so at the expense of weaker instruction following — eye-gaze inspections reveal this model to be struggling with instruction comprehension and adherence to formatting rules — leading to lower relative judgment performance. In contrast, Claude does not rank low as a judge despite its weak performance on U-MATH. Meanwhile, Gemini, known to excel in both mathematical problem solving and instruction following, comes out as the top-ranked judge.
Appendix I -MATH Behavior of Judges
In Figure 15 we visualize the difference in ‘performance profiles’ of the judges which we’ve discussed in Section 4.3 — proprietary models behaving more conservatively and Qwen family models exhibiting the opposite tendencies.
This is in line with eye-gazing inspections suggesting that:
Qwens tend to ‘follow the solution’ and are also good at going into involved derivation chains necessary to arrive at a true positive verdict in more complex scenarios, albeit at the cost of increased hallucination risk.
Proprietary models are more ‘anchored on the label’ and less heavy on long hallucination-prone transformation chains, which comes at the expense of missing more complex true positives.
Notably, Claude Sonnet 3.5 and Qwen2.5-Math 72B are the ‘opposite extremes’ of the observed patterns — having respectively the highest overall TNR and highest overall TPR — with an approximately equal F1-score. To illustrate the patterns, Appendix J provides an example comparing the Claude’s and Qwen’s judgments on a single -MATH sample. Notice how Claude is restrictive and superficial in its comparison, whereas Qwen ‘loses the structure’ along the way, designating only the first two steps prescribed with the CoT prompt (see prompt contents in Appendix C.2) and omitting points three and four.
These effects are further amplified with reduction in model size: proprietary models mainly lose in TPR when moving from larger models to smaller ones, whereas Qwens, once again on the contrary, lose more in TNR. A possible interpretation is that model size helps to mitigate the natural tendencies potentially induced by the training data, perhaps due to better generalization across general and domains-specific skills, or due to increased reliability.
Also, among the higher-ranking judges, the large Qwen2.5 and Gemini models turn out to be the two most balanced representatives of their respective source classes. We speculate that these two models are trained with data explicitly balancing math-specific skills and general capabilities: Gemini being heavily optimized for both math and instruction following, and Qwen likely trained on a blend containing synthetic data produced by various specialist models, including mathematical synthetic data from Qwen2-Math.
The behavioral differences may also be observed with predicted label agreement rates between judges, see Figure 16 for the comparison. Interestingly, no pair of models has agreement above around 80% — even for same-family models like Qwen2.5 and Qwen2.5-Math — despite the pairwise -MATH performance deltas being small compared to 20% disagreement.
All of this shows that judge comparison is substantive beyond the one-dimensional choice of the better model and suggests judge ensembling to be a potentially fruitful approach to evaluation.
Appendix J -MATH Judgment Examples
Appendix K Full -MATH Results
Model -MATH -MATH -MATH -MATH -MATH F1 TPR TNR PPV NPV F1 TPR TNR PPV NPV F1 TPR TNR PPV NPV F1 TPR TNR PPV NPV F1 TPR TNR PPV NPV Llama-3.1 8B 52.0 48.7 55.9 56.0 48.5 48.7 45.2 53.4 56.5 42.2 49.2 45.4 54.0 35.5 63.9 51.2 46.4 56.5 53.3 49.7 55.5 55.0 62.2 77.0 37.5 Qwen2.5 7B 69.3 78.7 59.8 69.3 70.8 62.4 75.5 49.1 66.5 60.0 72.3 78.4 70.1 59.4 85.3 68.3 81.4 55.7 66.3 73.7 69.1 79.4 59.8 82.0 55.7 Qwen2.5-Math 7B 61.9 76.6 47.9 62.9 63.9 59.7 71.0 48.3 64.7 55.4 63.8 85.6 51.7 49.7 86.5 57.2 75.0 41.2 57.7 60.7 63.8 77.8 50.0 78.2 49.4 GPT-4o-mini 72.3 59.0 88.1 85.1 65.1 69.3 56.1 87.1 85.3 59.8 76.2 59.8 90.2 77.3 80.1 70.4 56.4 86.3 81.4 64.9 69.6 63.0 87.8 92.2 50.7 Gemini 1.5 Flash 74.8 63.3 88.3 86.2 67.6 71.2 60.6 85.3 84.7 61.9 80.6 67.0 92.0 82.3 83.3 70.1 57.9 84.0 79.4 65.1 73.9 67.7 91.5 94.8 55.1 Llama-3.1-70B 61.0 62.5 59.6 64.1 57.9 56.0 58.7 53.4 62.8 49.2 57.0 63.9 54.0 43.7 72.9 69.4 67.1 71.8 71.8 67.1 58.8 61.4 61.0 78.4 40.7 Qwen2.5 72B 75.6 77.1 74.2 77.5 73.7 70.5 76.1 64.7 74.2 67.0 79.3 75.3 83.9 72.3 85.9 73.7 76.4 71.0 73.8 73.8 74.2 79.4 72.0 86.7 60.2 Qwen2.5-Math 72B 74.0 80.9 66.8 73.8 75.2 69.3 81.3 56.9 71.6 69.5 77.3 81.4 76.4 65.8 88.1 68.2 77.9 58.8 66.9 71.3 76.8 82.5 73.2 87.6 64.5 GPT-4o 77.4 70.1 85.9 85.1 71.3 74.2 67.1 83.6 84.6 65.5 81.8 71.1 90.8 81.2 84.9 77.5 72.1 83.2 82.1 73.6 72.6 70.4 82.9 90.5 54.8 Claude 3.5 Sonnet 74.8 62.5 89.5 87.3 67.4 70.8 58.1 87.9 86.5 61.1 77.9 59.8 93.1 82.9 80.6 72.2 57.9 88.5 84.4 66.3 73.8 70.9 85.4 91.8 56.0 Gemini 1.5 Pro 80.7 77.5 84.5 85.2 76.4 77.7 75.5 81.0 84.2 71.2 83.6 75.3 90.8 82.0 86.8 78.2 76.4 80.2 80.5 76.1 79.5 81.0 82.9 91.6 65.4
Model -MATH -MATH -MATH -MATH -MATH F1 TPR TNR PPV NPV F1 TPR TNR PPV NPV F1 TPR TNR PPV NPV F1 TPR TNR PPV NPV F1 TPR TNR PPV NPV Llama-3.1 8B 53.1 36.3 75.0 62.6 50.5 49.6 34.8 70.7 61.4 44.8 51.2 29.9 73.0 38.2 65.1 57.6 41.4 77.1 65.9 55.2 50.5 37.0 81.7 82.4 36.0 Qwen2.5 7B 67.0 65.9 68.4 70.7 63.5 60.5 61.3 60.3 67.4 53.8 72.4 71.1 75.3 61.6 82.4 66.4 67.1 65.6 67.6 65.2 65.0 66.1 69.5 83.3 47.1 Qwen2.5-Math 7B 61.2 72.8 49.9 62.7 61.4 56.7 67.7 45.7 62.5 51.5 64.0 80.4 55.2 50.0 83.5 58.5 68.6 48.9 58.9 59.3 61.2 76.2 46.3 76.6 45.8 GPT-4o-mini 69.2 50.8 92.0 88.1 61.8 61.7 43.2 87.9 82.7 53.7 78.5 58.8 94.8 86.4 80.5 69.8 50.7 92.4 87.7 63.7 64.3 52.9 91.5 93.5 45.7 Gemini 1.5 Flash 65.3 42.0 95.8 92.1 58.9 61.9 39.4 94.8 91.0 53.9 70.8 41.2 97.7 90.9 74.9 65.3 40.7 96.2 91.9 60.3 59.7 45.5 92.7 93.5 42.5 Llama-3.1-70B 68.2 60.9 76.7 75.2 63.0 63.8 58.7 70.7 72.8 56.2 70.2 60.8 79.3 62.1 78.4 69.8 58.6 82.4 78.1 65.1 64.4 64.6 70.7 83.6 46.4 Qwen2.5 72B 75.1 73.0 77.5 79.0 71.3 68.9 64.5 75.0 77.5 61.3 80.1 76.3 84.5 73.3 86.5 73.4 74.3 72.5 74.3 72.5 73.8 77.2 74.4 87.4 58.7 Qwen2.5-Math 72B 75.5 77.1 74.0 77.4 73.7 68.8 72.3 65.5 73.7 63.9 79.8 77.3 83.3 72.1 86.8 69.2 73.6 64.9 69.1 69.7 80.4 83.6 80.5 90.8 68.0 GPT-4o 74.2 61.4 89.3 86.9 66.7 68.2 56.8 83.6 82.2 59.1 78.9 61.9 93.1 83.3 81.4 75.8 65.0 87.8 85.0 70.1 70.5 62.4 91.5 94.4 51.4 Claude 3.5 Sonnet 68.1 46.8 95.2 91.9 60.8 64.1 43.2 94.0 90.5 55.3 71.8 41.2 99.4 97.6 75.2 68.1 45.7 95.4 91.4 62.2 63.4 53.4 87.8 91.0 45.0 Gemini 1.5 Pro 69.1 49.4 93.8 90.3 61.6 64.6 44.5 93.1 89.6 55.7 74.8 51.5 94.8 84.7 77.8 68.6 46.4 95.4 91.5 62.5 64.9 54.5 90.2 92.8 46.2

### v3

@@
 Our findings reveal that LLMs achieve a maximum accuracy of only 63% on text-based tasks, with even lower 45% on visual problems.
-The solution assessment proves challenging for LLMs, with the best LLM judge having an F1-score of 80% on -MATH
+The solution assessment proves challenging for LLMs, with the best LLM judge having an F1-score of 80% on -MATH.
 We open-source U-MATH, -MATH, and evaluation code on GitHub.111https://github.com/toloka/u-math
@@
 Reference solutions and answers accompany all problems.
-Accuracy is the primary performance metric for U-MATH, its text-only problems (U-MATH) and problems that include a visual component (U-MATH). The main performance measure for -MATH is macro-F1.
-We use an LLM as a judge (Zheng et al., 2023) to measure the accuracy of the free-form answers against the golden solutions.
+Accuracy is the primary performance metric for U-MATH, its text-only subset (U-MATH) and the subset of problems including a visual component (U-MATH).
+We use an LLM-as-a-judge (Zheng et al., 2023) to measure the accuracy of the free-form answers against the golden solutions.
 A problem is considered solved only if all required questions are answered and all requested items (e.g., all saddle points) are correctly identified.
@@
 The U-MATH benchmark comprises 1,100 carefully curated and validated mathematical problems.
-These problems are distributed across 6 core subjects with about 20% of the tasks incorporating visual elements, such as graphs, tables, and geometric figures, mirroring the multi-modal nature of real-world mathematical problems: Precalculus (Review), Algebra, Differential Calculus (+Differential Equations), Integral Calculus, Multivariable Calculus, and Sequences & Series.
+These problems are distributed across 6 core subjects with about 20% of the tasks incorporating visual elements, such as graphs, tables, and geometric figures, mirroring the multi-modal nature of real-world mathematical problems: Precalculus, Algebra, Differential Calculus (+Differential Equations), Integral Calculus, Multivariable Calculus, and Sequences & Series.
 Math Subject #Textual #Visual Avg.
@@
 Multivar C.
-Precalculus Seq.& Series T V T V T V T V T V* T V* T V* 900 200 150 30 150 70 150 58 150 28 150 10 150 4 Text-only models Mathstral 7B 18.0 20.7 6.0 51.3 6.7 4.0 10.0 1.3 1.7 8.0 3.6 48.7 10.0 10.7 0.0 NuminaMath 7B 19.2 22.8 3.0 62.7 0.0 4.0 7.1 1.3 0.0 6.0 3.6 51.3 0.0 11.3 0.0 Llama-3.1 8B 22.3 26.1 5.0 59.3 3.3 6.7 5.7 9.3 3.4 11.3 3.6 54.7 10.0 15.3 25.0 Qwen2.5 7B 33.8 40.0 6.0 86.0 10.0 12.7 1.4 10.0 12.1 26.7 3.6 75.3 0.0 29.3 0.0 Qwen2.5-Math 7B 38.4 45.2 7.5 87.3 6.7 18.7 5.7 8.0 10.3 36.0 10.7 80.7 0.0 40.7 0.0 NuminaMath 72B 25.0 29.7 4.0 74.7 3.3 6.7 4.3 4.0 3.4 11.3 3.6 62.7 10.0 18.7 0.0 Llama-3.1 70B 28.5 33.7 5.0 82.0 3.3 10.7 5.7 4.0 5.2 14.0 3.6 64.0 0.0 27.3 25.0 Llama-3.1 Nemotron 70B 31.4 37.4 4.0 84.0 0.0 14.7 2.9 4.0 3.4 25.3 7.1 64.0 20.0 32.7 0.0 Qwen2.5 72B 41.0 48.6 7.0 88.7 6.7 22.7 4.3 12.0 6.9 40.0 17.9 83.3 0.0 44.7 0.0 Athene-V2 72B Chat 46.2 54.6 8.5 88.7 3.3 34.0 4.3 16.0 6.9 50.7 21.4 88.7 10.0 49.3 50.0 Qwen2.5-Math 72B 50.2 59.0 10.5 92.7 6.7 35.3 7.1 20.7 17.2 58.0 7.1 90.0 0.0 57.3 50.0 Multimodal models Pixtral 12B 15.5 15.6 15.5 44.7 23.3 1.3 34.3 0.7 0.0 3.3 0.0 32.0 0.0 11.3 0.0 Llama-3.2 11B Vision 17.0 18.6 10.0 54.0 10.0 1.3 20.0 1.3 1.7 4.7 3.6 43.3 10.0 6.7 0.0 LLaVA-OV Qwen2-7B 17.7 20.7 4.5 60.7 6.7 4.0 5.7 1.3 1.7 5.3 3.6 43.3 10.0 9.3 0.0 Qwen2-VL 7B 20.4 21.4 15.5 62.7 10.0 4.7 32.9 0.7 5.2 6.7 7.1 45.3 0.0 8.7 0.0 Qwen2-VL 72B 31.2 32.2 26.5 80.7 26.7 9.3 40.0 2.0 13.8 14.7 28.6 65.3 10.0 21.3 0.0 Llama-3.2 90B Vision 32.6 36.3 16.0 85.3 26.7 10.7 25.7 2.7 1.7 22.7 7.1 65.3 20.0 31.3 25.0 Claude Sonnet 3.5 35.1 36.1 30.5 76.0 33.3 12.0 41.4 7.3 17.2 21.3 28.6 65.3 30.0 34.7 25.0 GPT-4o-mini 37.2 40.3 23.0 88.0 16.7 16.7 31.4 4.0 10.3 24.0 35.7 77.3 20.0 32.0 25.0 GPT-4o 43.5 46.4 30.0 91.3 30.0 18.7 32.9 10.0 20.7 41.3 42.9 79.3 30.0 38.0 25.0 Gemini 1.5 Flash 51.3 53.8 40.0 91.3 50.0 36.0 45.7 14.0 24.1 44.0 50.0 80.7 30.0 56.7 50.0 Gemini 1.5 Pro 60.1 63.4 45.0 91.3 60.0 50.7 47.1 27.3 24.1 60.7 57.1 87.3 70.0 63.3 50.0
+Precalculus Seq.& Series T V T V T V T V T V T V* T V* 900 225 150 30 150 80 150 68 150 28 150 13 150 6 Text-only models Mathstral-7B-v0.1 16.36 20.1 1.3 56.0 3.3 4.0 1.5 2.0 1.2 6.0 0.0 40.7 0.0 12.0 0.0 Llama-3.1-8B 15.47 19.0 1.3 52.0 3.3 4.7 2.9 1.3 0.0 10.0 0.0 36.7 0.0 9.3 0.0 Qwen2.5-7B 32.18 39.8 1.8 77.3 3.3 14.0 0.0 10.7 1.2 34.7 7.1 72.7 0.0 29.3 0.0 Qwen2-Math-7B 28.62 35.1 2.7 77.3 3.3 9.3 1.5 8.0 2.5 23.3 7.1 70.7 0.0 22.0 0.0 Qwen2.5-Math-7B 33.24 41.0 2.2 64.0 0.0 16.7 1.5 21.3 0.0 35.3 7.1 72.0 7.7 36.7 16.7 Llama-3.1-70B 26.13 32.0 2.7 70.0 10.0 10.0 0.0 5.3 2.5 25.3 3.6 55.3 0.0 26.0 0.0 Qwen2.5-72B 34.04 41.9 2.7 69.3 0.0 25.3 0.0 10.7 1.2 39.3 14.3 70.0 0.0 36.7 16.7 Qwen2-Math-72B 35.20 43.6 1.8 80.0 0.0 22.7 0.0 16.0 2.5 32.0 0.0 76.7 0.0 34.0 33.3 Qwen2.5-Math-72B 41.16 50.6 3.6 73.3 3.3 33.3 0.0 23.3 7.5 48.7 3.6 82.0 0.0 42.7 0.0 Multimodal models Qwen2-VL-7B 17.33 20.3 5.3 51.3 6.7 8.7 10.3 1.3 1.2 6.0 3.6 44.0 7.7 10.7 0.0 LLaVA-OV(Qwen2-7B) 14.40 17.8 0.9 48.7 0.0 4.7 1.5 0.7 0.0 7.3 3.6 36.7 0.0 8.7 0.0 Pixtral-12B-2409 15.64 18.1 5.8 46.7 16.7 4.7 10.3 0.7 1.2 6.7 0.0 41.3 0.0 8.7 0.0 Qwen2-VL-72B 26.93 30.1 14.2 70.0 13.3 11.3 20.6 5.3 7.5 18.7 21.4 58.7 15.4 16.7 0.0 GPT-4o 36.53 41.8 15.6 80.0 16.7 22.0 14.7 11.3 13.8 38.7 25.0 68.0 15.4 30.7 0.0 Gemini-1.5-Pro 48.89 53.4 30.7 84.7 56.7 37.3 27.9 27.3 22.5 42.7 28.6 81.3 46.2 47.3 16.7
 Among text-only models, the math-specific model Qwen2.5-Math-72B achieves the highest overall accuracy at 50.2%, showcasing strong mathematical reasoning capabilities.
@@
 Acknowledgement
+Some of the problems in U-MATH and mu-MATH are sourced from OpenStax under CC BY 4.0.
 We thank all contributors from Stevens Institute and Toloka.ai experts who assisted in sourcing and verifying problems, inspect the solutions and provided valuable feedback throughout the development of U-MATH.
-We would like to give special thanks to the dedicated team of Gradarius experts who played a crucial role in validating problems and ensuring their quality: Jan Cannizzo Paul Schwartz Andrey Nikolaev Arina Voorhaar Chloe Weiers Funda Gul Tatiana Zenkevich Igor Teplukhovskiy Anastasiia Feklina Ruslan Akmetdinov Ekaterina Eremina Sofia Tekhazheva
+We would like to give special thanks to the dedicated team of experts from Gradarius and Stevens Institute who played a crucial role in validating problems and ensuring their quality: Jan Cannizzo Paul Schwartz Andrey Nikolaev Arina Voorhaar Chloe Weiers Funda Gul Tatiana Zenkevich Igor Teplukhovskiy Anastasiia Feklina Ruslan Akmetdinov Ekaterina Eremina Sofia Tekhazheva
 Ethics Statement